# Optimal joint cutting of two-qubit rotation gates

### v1

@@ I Introduction
with the real coefficients $a_{i}$, to reduce the circuit size or the impact of noise. Circuit cutting can be categorized into two methods. Wire cutting [14, 15, 16, 17, 18, 19] describes the decomposition of the identity channel corresponding to empty qubit wires. Here, $\mathcal{F}_{i}$ are measure-and-prepare channels. The second method is known as gate cutting [20, 21, 22, 23, 24] where $\mathcal{F}_{i}$ are elements of LOCC($A$,$B$), that is local operations on two partitions $A$ and $B$ of the quantum circuit and classical communication between them. If all wires or gates that connect the partitions $A$ and $B$ are cut, the sub-circuits become independent and can be run on two quantum computers only connected by classical communication links. To evaluate Eq. 1, it can be rewritten as a quasi-probability decomposition (QPD) and evaluated via Monte-Carlo sampling [25]. As a result, all circuit-cutting methods incur a sampling overhead, the factor of more samples required to estimate the expectation value of an observable with the same accuracy as with respect to the original uncut circuit. This sampling overhead has been shown to be $\kappa^{2}$ where $\kappa=\sum_{i}|a_{i}|$ [25, 26]. The task is, therefore, to find a decomposition of $\mathcal{V}$ with minimal $\kappa$, which we will refer to as $\gamma$.
@@
Note the similarity of the approach to probabilistic error cancellation where the (possibly unphysical) inverse of a noisy quantum channel is decomposed into physical operations [27]. Recent work [23] employed CNOT-gate teleportation [28, 29, 30] consuming one Bell state initially connecting the two partitions per CNOT gate. The optimal QPD with $\gamma=3$ of the Bell-state density matrix then translates into the optimal decomposition of the gate. The crucial insight of Ref. [23] was that the joint QPD of $n$ Bell states required for the teleportation of $n$ CNOT gates can be constructed more efficiently compared to the individual decomposition of the gates. As a result, the $\gamma$ parameter is reduced from $\mathcal{O}(3^{n})$ to $\mathcal{O}(2^{n})$, however at the cost of one ancilla qubit per partition and gate. In this cutting scheme, $\mathcal{F}_{i}\in\mathrm{LOCC}(A,B)$ and two-way classical information is shared between the partitions. The protocol was extended to general Clifford gates and later also to wire cutting [15]. Recently, it was shown that optimal cutting of parallel wires is possible without ancilla qubits [16].
In this work, we extend previous proposals to the joint cutting of non-Clifford two-qubit rotation gates. We significantly simplify existing methods and, by that, enable implementation on current NISQ hardware. Two-qubit rotation gates play a crucial role for example in the QAOA algorithm [4] or for the simulation of spin systems. In Sec. II, we introduce a virtual teleportation scheme that effectively consumes less than one entanglement bit (ebit). After explaining the protocol for a single gate instance, we generalize the scheme to the joint cutting of $n$ two-qubit rotation gates and demonstrate the optimality of the derived circuit-cutting method. With optimality, we refer to the circuit cutting scheme with the minimal possible sampling overhead. The $\gamma$ parameter for a two-qubit rotation gate with parameter $\theta$ is [21, 23]
@@
asymptotically as $n$ approaches infinity, which matches the lower bound. Our scheme offers several improvements over previous work. First, it eliminates the need for classical communication between the partitions, and consequently, no real-time feedback is required. This contrasts the gate-teleportation protocol used in Ref. [23] where local correction operators conditioned on intermediate measurement results are necessary. Moreover, our protocol can be executed sequentially on the same quantum computer, unlike schemes where two-way classical communication exchange necessitates the use of two separate quantum computers in parallel. Second, it has recently been emphasized that the number of operations $\mathcal{F}_{i}$ can be the limiting factor for the execution of circuit-cutting protocols. We derive a quasi-probability decomposition of the entanglement resource states with an exponentially reduced number of elements. Third, in Sec. III, we show that parallel gates can be cut without the need for ancilla qubits. Parallel gates refer to gates that can be executed within the same time slice of a circuit as shown in Fig. 1. Note that controlled rotation gates and CNOT gates (for $\theta=\pi/2$) are equivalent to two-qubit rotation gates up to local operations. Thus, all results of our work also apply to these types of gates.
@@ II Virtual gate teleportation
In this section, we aim to find an optimal joint decomposition of two-qubit rotation gates based on gate teleportation. A teleportation scheme for a two-qubit rotation gate $R_{zz}(\theta)$ [30] is shown in Fig. 2. It consumes one initial Bell state (wiggly line in the figure) prepared on two ancilla qubits. These qubits are subsequently measured in the computational basis, and the outcomes, $k,l\in\{0,1\}$, are communicated between the partitions $A$ and $B$ via classical communication to select the local correction operations $X^{k}$ and $Z^{l}$. For $\theta=\pi/2$ the two-qubit rotation gate is equivalent to a CNOT gate up to local operations.
Figure 2: Teleportation protocol for a two-qubit rotation gate $R_{zz}(\theta)$ [30]. The scheme consumes an initial Bell state (wiggly line) as entanglement resource between the partitions $A$ and $B$ and requires two-way classical communication visualized by the dashed arrows. The measurements are performed in the computational basis with the outcomes $l,k\in\{0,1\}$.
@@ II.1 Single gate instance
Figure 3: Virtual gate teleportation. As explained in the main text, for the state shown in Eq. 4, the correct gate is only teleported for the measurement results $k=l$, resulting in suboptimal sampling overhead when discarding other outcomes. Since we are only interested in a teleportation scheme together with a quasi-probability simulation of the initial state, we show that in this case, the result $k\neq l$ can be corrected in the final classical post-processing step. This effectively corresponds to the preparation of the initial state $|\Psi_{kl}\rangle$ defined in Eq. 10, which depends on the outcome of the measurement performed later in time, therefore the term virtual teleportation.
@@
in which the signs of the quasi-probabilities depend on the measurement outcomes in the form of the factor $(-1)^{k+l}$. To execute the protocol, we proceed as detailed in the following: To estimate an expectation value experimentally, one typically defines a post-processing function $f(s)\in[-1,1]$ [14] as a function of the bitstring $s$ observed at each experimental run and then takes the sample mean over many runs. For our protocol, we perform Monte-Carlo sampling as, for example, explained in Ref. [24]. For each run, we prepare one of the states in Eq. 8 and Eq. 9 sampled according to the probability given by the absolute of the prefactor divided by $\gamma$. In case we initialize the ancilla qubits with one of the states constituting $\sigma^{+}$ or $\sigma^{-}$, we accumulate ±$\pm(-1)^{k+l}f(s)$ in the sample mean where the sign depends on the outcome of the measurements on the ancilla qubits. We will refer to our protocol as virtual gate teleportation since an actual entanglement resource state is not physically prepared but only simulated by a quasi-probability decomposition. In this respect, the protocol effectively makes use of the state
@@ II.2 Joint cutting of multiple gates
Note that no classical communication is required between the partitions  [^1] It can be shown that all operations in the virtual teleportation protocol are elements of the extended definition of local operations used in Ref. [23, 15].. Joint gate cutting without the need for classical communication significantly alleviates implementation on current hardware for two reasons. First, no real-time feedback for local operations conditioned on measurement results is needed, distinct from conventional gate teleportation protocols as, for example, shown in Fig. 2. Second, if all gates connecting the two partitions are cut, the severed sub-circuits can be executed sequentially on the same hardware. This stands in stark contrast to protocols employing two-way classical communications, where two quantum devices connected by classical links must operate in parallel. Lastly, we mention again that our protocol contains the joint cutting of $n$ CNOT gates as a special case with optimal $\gamma_{\mathrm{joint}}^{(n)}=2^{n+1}-1$ [23].
@@ III Parallel gates can be cut without ancilla qubits
Interestingly, as we will show in the following, there is an ancilla-free joint decomposition of parallel gates. With parallel gates, we refer to gates that can be executed at the same time slice of the circuit, for example, as shown in Fig. 1 for the case of CNOT gates. In Appendix C, we start from the protocol in Fig. 4 and subsequently show that it can be reduced to an ancilla-free decomposition in terms of local operations on the partition $A$ and $B$. Remarkably, again, no classical communication between the partitions is required, and optimality is preserved. The explicit formula for the decomposition is shown in Eq. 42. It contains single-qubit $Z$ gates, as well as multi-qubit operations on all subsets of the $n$ qubits on each partition $A$ and $B$. The operations consist of two different types. The first are multi-qubit rotation gates by angles ±$\pm\pi/2$. They can be implemented by single-qubit rotation gates sandwiched by ladders of CNOT gates [38]. The second are non-unitary operations, where the single-qubit rotation gates within the CNOT ladders are replaced by computational-basis measurements; see a detailed discussion in Appendix C. As a consequence, each term in the decomposition can be implemented with $\mathcal{O}(n)$ CNOT gates. It has recently been emphasized that the sampling overhead is not the only relevant metric measuring the quality of a circuit-cutting scheme [16] but also the number $L$ of different channels to be applied. In principle, the number of experimental shots required to evaluate an expectation value to additive error $\epsilon$ scales as $\mathcal{O}(\gamma^{2}/\epsilon^{2})$ which is independent of $L$. In practice, however, the latency introduced by the additional compilation time needed for each of the $L$ sub-circuits corresponding to the elements $\mathcal{F}_{i}$ for $i=1,...,L$ on topologically constrained hardware can be the dominant factor in the runtime of the algorithm [16]. As discussed in Appendix A, our decomposition guarantees $L=\mathcal{O}(4^{n})$ for the case of parallel as well as non-parallel gates, which is an exponential improvement over previous results. It should be noted that we could trade compilation time versus additional circuit depth. Indeed, a multi-qubit rotation gate can be trivially routed to topologically constrained hardware by SWAP insertion. This decreases compilation latency to almost zero but, in turn, might drastically increase the CNOT-gate count and the additional circuit depth $d$, resulting in a more severe impact of noise. Note that a multi-qubit rotation gate can be implemented on fully connected hardware with depth $d=\mathcal{O}(\lceil\mathrm{log}(n)\rceil)$ when replacing each CNOT gate ladder by a balanced tree [38]. For hardware with constrained connections, efficient compilation methods exist for multi-qubit rotation gates based on minimum spanning trees [39] and CNOT gate re-synthesis [40]. In what sense the low depth of the circuits can be retained will be the focus of further investigations.
@@ Appendix C Parallel gates
This state is just

### v2

@@ I Introduction
with the real coefficients $a_{i}$, to reduce the circuit size or the impact of noise.
Circuit cutting can be categorized into two methods. Wire cutting [14, 15, 16, 17, 18, 19] effectively corresponds to the cutting of horizontal empty qubit wires which are modeled mathematically by the identity channel. Consequently, wire cutting is a decomposition of this channel into measure-and-prepare channels. The second method, on which the focus will be in this work, is known as gate cutting [20, 21, 22, 23, 24]. Here, $\mathcal{F}_{i}$ are elements of LOCC($A$,$B$), that is local operations on two partitions $A$ and $B$ of the quantum circuit and classical communication between them. Since $\mathcal{F}_{i}$ are local operations, no entanglement is created by the channel, but is effectively simulated as described below. If all wires or gates that connect the partitions $A$ and $B$ are cut, the sub-circuits become independent and can be run on two quantum computers only connected by classical communication links. Most circuit-cutting research so far has focused on applications to algorithms where the output is an expectation value of an observable. To evaluate Eq. 1 in this context, we first define $\kappa=\sum_{i}|a_{i}|$ and the probability distribution $p_{i}=|a_{i}|/\kappa$. Next, Eq. 1 is rewritten as
$\mathcal{V}=\kappa\sum_{i}p_{i}\mathrm{sign}(a_{i})\mathcal{F}_{i}$ (2)
@@
and evaluated via Monte-Carlo sampling. At each experimental shot we select the $i$th channel $\mathcal{F}_{i}$ with probability $p_{i}$ and evaluate the circuit with $\mathcal{V}$ replaced by $\mathcal{F}_{i}$. The measurement outcome is weighted by $\kappa\mathrm{sign}(a_{i})$ and the mean over many runs produces an unbiased estimate for the expectation value of the observable. This sampling procedure is referred to as quasi-probability sampling [25] because of the appearance of the sign of $a_{i}$ in Eq. 2 which is referred to as a quasi-probability decomposition (QPD).
Quasi-probability sampling incurs a sampling overhead, the factor of more samples required to estimate the expectation value of an observable with the same accuracy as with respect to the original uncut circuit. This sampling overhead has been shown to be $\kappa^{2}$ [25, 26] and originates from the factor of $\kappa$ in front of the sum in Eq. 2.
Note the similarity of the approach to probabilistic error cancellation where the (possibly unphysical) inverse of a noisy quantum channel is decomposed into physical operations [27].
The application of circuit cutting to sampling tasks has been studied initially in Ref. [17] and recently more extensively in Ref. [28] with the result that circuit cutting can also be meaningful in this situation.
When gates or wires are cut individually, the overall sampling overhead increases exponentially in the number of gates and wires cut. This strong increase of sampling overhead renders circuit cutting prohibitively expensive when the cutting location in a circuit and the cutting scheme is not carefully selected. The task is, therefore, to find decompositions of gates or collections of gates with minimal $\kappa$, which we will refer to as $\gamma$.
Ref. [23] employed CNOT-gate teleportation [29, 30, 31] consuming one Bell state initially connecting the two partitions per CNOT gate. The optimal QPD with $\gamma=3$ of the Bell-state density matrix then translates into the optimal decomposition of the gate. The crucial insight of Ref. [23] was that the joint QPD of $n$ Bell states required for the teleportation of $n$ CNOT gates can be constructed more efficiently compared to the individual decomposition of the gates. As a result, the $\gamma$ parameter is reduced from $\mathcal{O}(3^{n})$ to $\mathcal{O}(2^{n})$, however at the cost of one ancilla qubit per partition and gate. In this cutting scheme, $\mathcal{F}_{i}\in\mathrm{LOCC}(A,B)$ and two-way classical information is shared between the partitions. The protocol was extended to general Clifford gates and later also to wire cutting [15]. Recently, it was shown that optimal cutting of parallel wires is possible without ancilla qubits [16].
In this work, we extend previous proposals to the joint cutting of non-Clifford two-qubit rotation gates. We significantly simplify existing methods and, by that, enable implementation on current noisy intermediate-scale quantum (NISQ) hardware. Two-qubit rotation gates play a crucial role for example in the quantum approximate optimization algorithm (QAOA) [4] or for the simulation of spin systems. In Sec. II, we introduce a virtual teleportation scheme that effectively consumes less than one entanglement bit (ebit). After explaining the protocol for a single gate instance, we generalize the scheme to the joint cutting of $n$ two-qubit rotation gates and demonstrate the optimality of the derived circuit-cutting method. With optimality, we refer to the circuit cutting scheme based on quasi-probability sampling with the minimal possible sampling overhead. A two-qubit rotation gate with rotation angle $\theta$ is defined as
$R_{zz}(\theta)=\mathrm{cos}(\theta/2)\mathbb{I}\otimes\mathbb{I}-i\mathrm{sin}(\theta/2)Z\otimes Z$ (3)
with the single-qubit identity $\mathbb{I}$ and the Pauli $Z$ matrix. The $\gamma$ parameter for this gate is [21, 23]
@@
asymptotically as $n$ approaches infinity, which matches the lower bound. Our scheme offers several improvements over previous work. First, it eliminates the need for classical communication between the partitions, and consequently, no real-time feedback is required. This contrasts the gate-teleportation protocol used in Ref. [23] where local correction operators conditioned on intermediate measurement results are necessary. Moreover, as explained in more detail in Sec. II.B our protocol can be executed sequentially on the same quantum computer, unlike schemes where two-way classical communication exchange necessitates the use of two separate quantum computers in parallel. Second, it has recently been emphasized that the number of operations $\mathcal{F}_{i}$ can be the limiting factor for the execution of circuit-cutting protocols [16]. We derive a quasi-probability decomposition of the entanglement resource states with an exponentially reduced number of elements. Third, in Sec. III, we show that parallel gates can be cut without the need for ancilla qubits. Parallel gates refer to gates that can be executed within the same time slice of a circuit as shown in Fig. 1. Note that controlled rotation gates and CNOT gates (for $\theta=\pi/2$) are equivalent to two-qubit rotation gates up to local operations. Thus, all results of our work also apply to these types of gates.
@@ II Virtual gate teleportation
In this section, we aim to find an optimal joint decomposition of two-qubit rotation gates based on gate teleportation. A teleportation scheme for a two-qubit rotation gate $R_{zz}(\theta)$ [31] is shown in Fig. 2. It consumes one initial Bell state (wiggly line in the figure) prepared on two ancilla qubits. These qubits are subsequently measured in the computational basis, and the outcomes, $k,l\in\{0,1\}$, are communicated between the partitions $A$ and $B$ via classical communication to select the local Pauli correction operators $X^{k}$ and $Z^{l}$. For $\theta=\pi/2$ the two-qubit rotation gate is equivalent to a CNOT gate up to local operations.
Figure 2: Teleportation protocol for a two-qubit rotation gate $R_{zz}(\theta)$ [31]. The scheme consumes an initial Bell state (wiggly line) as entanglement resource between the partitions $A$ and $B$ and requires two-way classical communication visualized by the dashed arrows. The measurements are performed in the computational basis with the outcomes $l,k\in\{0,1\}$. Here, $H$ is the Hadamard gate and $Z$ and $X$ are the Pauli operators
@@ II.1 Single gate instance
Figure 3: Virtual gate teleportation. As explained in the main text, for the state shown in Eq. 6, the correct gate is only teleported for the measurement results $k=l$, resulting in suboptimal sampling overhead when discarding other outcomes. Since we are only interested in a teleportation scheme together with a quasi-probability simulation of the initial state, we show that in this case, the result $k\neq l$ can be corrected in the final classical post-processing step. This effectively corresponds to the preparation of the initial state $|\Psi_{kl}\rangle$ defined in Eq. 12, which depends on the outcome of the measurement performed later in time, therefore the term virtual teleportation. In the figure, $S$ is the phase gate.
@@
in which the signs of the quasi-probabilities depend on the measurement outcomes in the form of the factor $(-1)^{k+l}$. To execute the protocol, we proceed as detailed in the following: To estimate an expectation value experimentally, one typically defines a post-processing function $f(s)\in[-1,1]$ [14] as a function of the bitstring $s$ observed at each experimental run and then takes the sample mean over many runs. As explained in the introduction, we perform Monte-Carlo sampling as, for example, explained in Ref. [24]. For each run, we prepare one of the states in Eq. 10 and Eq. 11 sampled according to the probability given by the absolute of the prefactor divided by $\gamma$. In case we initialize the ancilla qubits with one of the states constituting $\sigma^{+}$ or $\sigma^{-}$, we accumulate ±$\pm(-1)^{k+l}\gamma f(s)$ in the sample mean where the sign depends on the outcome of the measurements on the ancilla qubits. We will refer to our protocol as virtual gate teleportation since an actual entanglement resource state is not physically prepared but only simulated by a quasi-probability decomposition. In this respect, the protocol effectively makes use of the state
@@ II.2 Joint cutting of multiple gates
Note that no classical communication is required between the partitions  [^1] It can be shown that all operations in the virtual teleportation protocol are elements of the extended definition of local operations used in Ref. [23, 15].. Joint gate cutting without the need for classical communication significantly alleviates implementation on current hardware for two reasons. First, no real-time feedback for local operations conditioned on measurement results is needed, distinct from conventional gate teleportation protocols as, for example, shown in Fig. 2. Second, if all gates connecting the two partitions are cut, the severed sub-circuits can be executed sequentially on the same hardware. Indeed, given a decomposition as in Eq. 1, we first determine the number of shots $N$ required to achieve a certain accuracy of the expectation value estimate with high probability. Next, we determine how many times the circuit with channel $\mathcal{F}_{i}$ needs to be executed by sampling $N$ times from the probability distribution with $p_{i}=|a_{i}|/\kappa$ and counting the number of times $i$ is realized. If the decomposition does not require classical communication, $\mathcal{F}_{i}=\mathcal{F}_{i}^{A}\otimes\mathcal{F}_{i}^{B}$ and if $\mathcal{F}_{i}$ on one partition involves measurements, the $i$th channel on the other partition is independent of the measurement result. If all gates connecting the two partitions are cut, the circuits can be evaluated sequentially and independently from each other since $\mathcal{F}_{i}^{A}$ and $\mathcal{F}_{i}^{B}$ are completely independent. After all circuits have been executed, the bitstrings that were measured on both partitions for the $i$th channel are concatenated and passed to the post-processing function. This procedure stands in stark contrast to protocols employing two-way classical communications where two quantum devices connected by classical links must operate in parallel. This results from the structure of $\mathcal{F}_{i}$ when allowing classical communication. Here, we allow operations that may depend on the outcome of a measurement on the other partition. Since measurement results are probabilistic, the number of times each channel has to be applied cannot be calculated prior to the experiment.
Lastly, we mention again that our protocol contains the joint cutting of $n$ CNOT gates as a special case with optimal $\gamma_{\mathrm{joint}}^{(n)}=2^{n+1}-1$ [23].
@@ III Parallel gates can be cut without ancilla qubits
Interestingly, as we will show in the following, there is an ancilla-free joint decomposition of parallel gates. With parallel gates, we refer to gates that can be executed at the same time slice of the circuit, for example, as shown in Fig. 1 for the case of CNOT gates. In Appendix C, we start from the protocol in Fig. 4 and subsequently show that it can be reduced to an ancilla-free decomposition in terms of local operations on the partition $A$ and $B$. Remarkably, again, no classical communication between the partitions is required, and optimality is preserved. The explicit formula for the decomposition is shown in Eq. 44. It contains single-qubit $Z$ gates, as well as multi-qubit operations on all subsets of the $n$ qubits on each partition $A$ and $B$. The operations consist of two different types. The first are multi-qubit rotation gates by angles ±$\pm\pi/2$. They can be implemented by single-qubit rotation gates sandwiched by ladders of CNOT gates [39]. The second are non-unitary operations, where the single-qubit rotation gates within the CNOT ladders are replaced by computational-basis measurements; see a detailed discussion in Appendix C. As a consequence, each term in the decomposition can be implemented with $\mathcal{O}(n)$ CNOT gates. Note that an $n$ qubit rotation gate about ±$\pm\pi/2$ can be implemented with exactly $n$ CNOT gates. Because of that and since the majority of terms in the decomposition of an $n$-qubit rotation gate shown in Eq. 44 contain operations on less than $n$ qubits, we expect strong noise reduction by using our scheme. Noise mitigation due to the reduction of costly two-qubit gates in circuit-cutting schemes was recently observed experimentally in Ref. [24].
It has been emphasized that the sampling overhead is not the only relevant metric measuring the quality of a circuit-cutting scheme [16] but also the number $L$ of different channels to be applied. In principle, the number of experimental shots required to evaluate an expectation value to additive error $\epsilon$ scales as $\mathcal{O}(\gamma^{2}/\epsilon^{2})$ which is independent of $L$. In practice, however, the latency introduced by the additional compilation time needed for each of the $L$ sub-circuits corresponding to the elements $\mathcal{F}_{i}$ for $i=1,...,L$ on topologically constrained hardware can be the dominant factor in the runtime of the algorithm [16]. As discussed in Appendix A, our decomposition guarantees $L=\mathcal{O}(4^{n})$ for the case of parallel as well as non-parallel gates, which is an exponential improvement over previous results. It should be noted that we could trade compilation time versus additional circuit depth. Indeed, a multi-qubit rotation gate can be trivially routed to topologically constrained hardware by SWAP insertion. This decreases compilation latency to almost zero but, in turn, might drastically increase the CNOT-gate count and the additional circuit depth $d$, resulting in a more severe impact of noise. Note that a multi-qubit rotation gate can be implemented on fully connected hardware with depth $d=\mathcal{O}(\lceil\mathrm{log}(n)\rceil)$ when replacing each CNOT gate ladder by a balanced tree [39]. For hardware with constrained connections, efficient compilation methods exist for multi-qubit rotation gates based on minimum spanning trees [40] and CNOT gate re-synthesis [41]. In what sense the low depth of the circuits can be retained will be the focus of further investigations.
@@ IV Conclusion
After completion of our manuscript, Refs. [42, 43] appeared as preprints which generalize some of our ideas.
@@ Appendix C Parallel gates
With the Hadamard gate $H$ and the phase gate $S$ this state is